Rationalize the denominator and simplify 5 divided by 4-√3

Answer by Boreal(15235)

Multiply top and bottom by the conjugate of the denominator
that would be 4 + sqrt(3)
the numerator becomes 5(4+ sqrt(3))=20+5 sqrt(3)
the denominator becomes (4- sqrt(3))(4+ sqrt(3))=16-3=13, because the internal products 4 sqrt(3) and -4 sqrt (3) cancel each other.
The top can be left as [5(4+sqrt (3)]/13